# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### 26 Mar 2026 v2

@@ Abstract
We present SciCoQA, a dataset for detecting discrepancies between scientific publications and their codebases to ensure faithful implementations. We construct SciCoQA from GitHub issues and reproducibility papers, and to scale our dataset, we propose a synthetic data generation method for constructing paper-code discrepancies. We analyze the paper-code discrepancies in detail and propose discrepancy types and categories to better understand the occurring mismatches. In total, our dataset consists of $635$ paper-code discrepancies ($92$ real, $543$ synthetic), covering the AI domain from real-world data and extending to Physics, Quantitative Biology, and other computational sciences through synthetic data. Our evaluation of $22$ LLMs demonstrates the difficulty of SciCoQA, particularly for instances involving omitted paper details, long-context inputs, and data outside the models’ pre-training corpus. The best-performing models in our evaluation, Gemini 3.1 Pro and GPT-5 Mini, detect only $46.7\%$ of real-world paper-code discrepancies.
@@ 1 Introduction
The “reproducibility crisis” in AI and across science casts doubt on the reliability of research (Baker, 2016; Hutson, 2018). To address this, the computational sciences have long recognized that a paper alone is insufficient and that publishing code, data, and instructions is a prerequisite to ensure experimental findings are reproducible (Buckheit and Donoho, 1995; Peng, 2011; Pineau et al., 2021).
@@
Figure 1: Example from SciCoQA, showing specific model implementation in the paper, and its implementation in the code (simplified for readability). The paper’s description and the code’s implementation mismatch, creating a paper-code discrepancy.
@@
However, the availability of code does not guarantee reproducibility nor consistency with the scientific text (as showcased in Fig. 1). In practice, implementation details can diverge from their descriptions, introducing performance variations that go unreported (Henderson et al., 2018). These discrepancies manifest in troubling ways: from “mathiness,” where equations simulate technical depth while actual gains stem from undocumented tricks (Lipton and Steinhardt, 2019), to evaluation metrics that differ in implementation (Post, 2018; Marie et al., 2021), rendering scientific comparisons invalid.
@@
Reproducibility issues and paper-code inconsistencies are typically only detected during reproduction efforts and post-publication, resulting in a waste of resources and eroding trust in science. While checklists during peer review can raise awareness of authors to provide (more) reproducible code (Pineau et al., 2021; Dodge et al., 2019), ideally, reviewers check the correctness of the implementation and ensure reproducibility. However, with the rapid growth of submission numbers, reviewers are already under severe time pressure (Rogers and Augenstein, 2020) and conducting intricate code reviews is time-consuming (Collberg and Proebsting, 2016).
@@
Furthermore, the reliance on manual review is becoming increasingly impractical as science begins to scale via agentic systems which develop ideas, generate and execute code, and produce scientific articles autonomously (Boiko et al., 2023; Lu et al., 2024; Tang et al., 2025; Weng et al., 2025b), a trend recently validated by peer-reviewed acceptance to an ICLR workshop (Yamada et al., 2025). While this direction has the potential to accelerate scientific discovery, it makes oversight increasingly challenging as humans cannot verify the rapidly expanding volume of output (Bowman et al., 2022). Moreover, the faithfulness of these systems is not guaranteed as LLMs suffer from limited context (Liu et al., 2024b), experience compounding errors (Dziri et al., 2023; Mirzadeh et al., 2025), and struggle with independent self-correction (Tyen et al., 2024; Wu et al., 2024).
@@
Consequently, ensuring the reliability of science at this scale demands automated tools capable of verifying that the code faithfully implements the methods reported in the scientific paper. To this end, we introduce the SciCoQA dataset for a novel, real-world quality assurance task: detecting discrepancies between scientific papers and code in the computational sciences. Benchmarking LLMs on this task allows us to measure their potential for automated quality assurance workflows in science. To construct our dataset, we leverage issues reported in code repositories and papers dedicated to reproducing existing work, such as those from reproducibility challenges and conference tracks. While these sources are realistic and of high quality, they are sparse and limited to CS and AI. Therefore, we scale the data to various computational science domains (e.g., Physics and Quantitative Biology) by generating modifications in codebases, creating discrepancies between the paper and its implementation. Finally, we benchmark open-weight and proprietary LLMs to determine whether current systems can be deployed for detecting paper-code discrepancies. Our evaluation reveals that the best-performing models are precise in identifying discrepancies; however, their recall is too low to guarantee that the code faithfully implements the paper, highlighting the need to develop more effective models and systems for this critical task.
@@ 2 Related Work
Recently, LLMs have been employed to detect errors in scientific papers. In an early work, Liu and Shah (2023) created short CS papers with logical errors and prompted LLMs to identify the issues. They found that while models could produce fluent reviews, they struggled to identify these fundamental errors without explicit instructions in the prompt. However, their reliance on manually crafted papers limits scalability. Therefore, Dycke and Gurevych (2025) and Xi et al. (2025) propose dedicated pipelines that generate modifications to papers that invalidate their key claims and evaluate whether the soundness score of generated reviews decreases or whether LLMs can spot these errors. Both find that LLMs struggle to detect most introduced errors. Besides constructing erroneous papers, recent research has also focused on identifying real-world issues. Specifically, Son et al. (2025) and Zhang and Abernethy (2025) use the authors’ retraction notes from WithdrarXiv (Rao et al., 2024) and PubPeer to obtain data of errors in scientific publication. While realistic, these retraction notes are often unspecific or describe errors on a high level (e.g., “Theorem 2 is incorrect”). Recently, Bianchi et al. (2025) developed a system to detect inconsistencies within the text of AI papers (e.g., incorrect calculations in tables, errors in equations, and imprecise definitions) and then manually verify the detections. They found their system to be relatively precise, but struggled with recall, i.e., detecting all issues in a paper. They also noted an increasing trend of errors in scientific publications over the years, highlighting the importance of quality assurance in science.
@@
While these works share a similar spirit to ours, i.e., identifying errors in science, they are limited by solely analyzing the paper. SciCoQA extends beyond the scientific text and considers the code as well, enabling the detection of cross-modal error types, such as implementations that deviate from the paper description or omissions in the paper or code that are crucial for understanding and reproducibility. Furthermore, SciCoQA utilizes constructed errors to scale the quantity and diversity in the computational science domain, while also providing challenging real-world discrepancies between paper and code.
@@
Our task is related to Code-Comment Inconsistency (CCI) detection, where natural language comments that no longer match a piece of code need to be identified, e.g., after adding a feature to a function. Early CCI research relied on rule-based systems to identify outdated comments (Ratol and Robillard, 2017). Subsequently, bi-encoders and cross-encoders have been employed (Rabbi and Siddik, 2020; Steiner and Zhang, 2022). Panthaplackel et al. (2021) further extended CCI to dynamic settings, learning to detect inconsistency arising from code updates. Recently, LLMs have been deployed for both detection and rectification (Dau et al., 2024; Rong et al., 2025). These methods deal with local and small inputs, typically taking a single function and an inline comment or docstring as input to detect the CCI. SciCoQA is crucially much more ambitious: it requires performing a global alignment, by reasoning over a dense scientific paper and a long, multi-file code repository to find inconsistencies.
@@
Automation in science is rapidly accelerating (Zheng et al., 2025). Early work targeted specific components of the research cycle, including ideation (Si et al., 2025; Baek et al., 2025), literature review Wang et al. (2024); Liang et al. (2025), coding (Tian et al., 2024; Gandhi et al., 2025), and writing (Liang et al., 2024; Weng et al., 2025a). A recent trend involves generating entire code repositories from a paper (Starace et al., 2025; Seo et al., 2025; Xiang et al., 2025) or automating the scientific method end-to-end, giving rise to “AI Scientists.” However, for both these tasks, validating whether the generated code faithfully reflects the paper is challenging. For example, PaperBench (Starace et al., 2025) relies on high-quality, but expensive, manual rubrics specific to each paper to verify implementation nuances. Others use generic LLM-judges without ground truth validation Tang et al. (2025), evaluate on whether the code runs (Weng et al., 2025b; Xiang et al., 2025), or assign reproducibility scores based on execution outcomes Siegel et al. (2024); Hu et al. (2025); Ye et al. (2025). Critically, these metrics can be misleading; a generated codebase may execute perfectly and achieve high performance while implementing a method that differs fundamentally from the paper’s scientific description (Beel et al., 2025). SciCoQA addresses this gap by providing a ground truth dataset of paper-code discrepancies, enabling the development of robust quality assurance models that can verify the faithfulness of (generated) papers to (generated) code. Thus, SciCoQA takes a step toward verification-oriented and reproducibility tooling essential for sustaining peer review integrity and science at scale (Wei et al., 2025b; Woodruff et al., 2026; You et al., 2026).
@@ 3 SciCoQA
The paper-code discrepancies, according to our definition, can occur from distinct authorship (where the paper writer differs from the engineer implementing the code), simplifications made in the text for readability, or code updates or experiments, that were not propagated to the manuscript.
@@ 3.1 Data Collection
Fig. 2 provides an overview of our data collection. To obtain real-world instances of paper–code discrepancies, we draw from GitHub issues and reproducibility papers. Furthermore, synthetic discrepancies are generated in real scientific codebases.
@@
We first identify repositories that reference a research paper in their homepage or description field on GitHub, restricted to projects published between 2020 and 2025 and repositories with at least one issue (see §A.1 for details). [^1] The crawled repositories are not always the official implementation by the authors, but sometimes reproductions. They were published, for example, when there is no official implementation or the reimplementation uses a different framework. We still consider these repositories as researchers might use them to base their experiments on. For each, we crawl all associated issues, yielding 1,8901{,}890 repositories with a total of 10,63610{,}636 issues. To process this large volume efficiently, we automatically classify issues with Qwen3 4B Thinking (Yang et al., 2025), prompting the model to determine whether an issue reports a paper–code discrepancy or not, resulting in $232$ candidates (prompt in §F.1). All candidates are subsequently manually filtered to ensure they meet our definition of a paper-code discrepancy, yielding $59$ discrepancies.
@@
We collect reproducibility reports from the ML Reproducibility Challenge (Pineau et al., 2019; Sinha et al., 2020, 2022, 2023) which invites participants to reproduce papers from leading ML, CV, and NLP conferences (e.g., NeurIPS, ICML, ICLR, CVPR, ICCV, ACL, EMNLP) and papers from the reproducibility tracks of SIGIR and ECIR from 2020–2025. We retain only reports that reproduce a single, open-access paper with an open-source implementation, resulting in $171$ reproducibility papers. From each report, we extract mentions of paper–code discrepancies using GPT-5 resulting in $132$ candidates (prompt in §F.3). Each extracted discrepancy is manually verified to confirm adherence to our definition, of which $65$ withstand our filter.
@@
Finally, all manually filtered discrepancies are validated by Gemini 3.1 Pro and GPT-5, checking the discrepancy’s existence given the original paper, codebase, and the GitHub issue or the description extracted from the reproducibility paper. For the cases where the LLMs disagreed, we manually verified the discrepancy. During this step, the LLMs are further tasked with generating a standardized discrepancy description. We prompt the model to output 3-8 sentences and state what the paper describes, what the code implements, and where the difference is. We use these generated descriptions as the ground truth to ensure the same format and level of verbosity for paper-code discrepancies (prompt in §F.2 and §F.4). We use two state-of-the-art LLMs to perform the validation and rephrasing to prevent potential bias in the selection and phrasing (see §C.5 for an analysis). The verification step ensures adherence to the discrepancy definition, acting as a high-precision filter. By retaining only discrepancies that can be verified given the explicit evidence, we prioritize the precision and objectivity of the ground truth. In total, we obtain $92$ discrepancies from $72$ papers, where $42$ discrepancies originate from GitHub issues and $50$ from reproducibility papers.
@@
To scale the data in size and beyond the CS/AI domain, we employ synthetic data generation. Specifically, based on our GitHub crawl, we sample repositories linked to arXiv papers and those with permissive code licenses (i.e., MIT, Apache 2.0, CC-BY, BSD), as we will be redistributing their code partially. We randomly select $102$ repositories where the paper’s arXiv subject is CS (balanced over Machine Learning (ML), cs.LG, Computer Vision (CV), cs.CV, and Natural Language Processing (NLP), cs.CL), and $102$ for non-CS papers. [^2] We manually checked the repositories and removed codebases from our sample that were not suitable, such as repositories for survey papers or with very few files. Next, we prompt GPT-5 with the paper and code to generate five code diffs according to our discrepancy definition (prompt in §F.5). We then sample up to three of these changes, ensuring that they do not manipulate the same file and that the generated diff can be found by exact match in the original code. Finally, we add $13$ additional discrepancies that are labeled as paper omissions as we later find these to be most critical (see §3.2 and §5). This yields a total of $543$ discrepancies across the $204$ papers, of which $286$ discrepancies are from the CS domain, and $257$ are from other computational domains, such as Electrical Engineering and Systems Science ($114$), Physics ($51$), or Statistics ($40$).

### 22 Apr 2026 v3

@@ Abstract
Discrepancies between scientific papers and their code undermine reproducibility, a concern that grows as automated research agents scale scientific output beyond human review capacity. Whether LLMs can reliably detect such discrepancies has not been systematically measured. To this end, we present SciCoQA, a dataset of $635$ paper-code discrepancies ($92$ real, $543$ synthetic) for this cross-modal verification task. Across $22$ evaluated models, even the best-performing LLMs, Gemini 3.1 Pro and GPT-5 Mini, detect only $46.7\%$ of real-world discrepancies, revealing a critical gap in automated scientific quality assurance. We construct SciCoQA from GitHub issues and reproducibility papers, and propose a synthetic generation pipeline to scale beyond AI to Physics, Quantitative Biology, and other computational sciences. We further introduce a taxonomy of discrepancy types and categories to characterize the occurring mismatches. Our analysis shows that models particularly struggle with omitted paper details, long-context inputs, and papers outside their pre-training corpus.  [^1] Data also available at: https://tudatalib.ulb.tu-darmstadt.de/handle/tudatalib/4994
@@ 1 Introduction
The “reproducibility crisis” in AI and across science casts doubt on the reliability of research (Baker, 2016; Hutson, 2018). To address this, the computational sciences have long recognized that a paper alone is insufficient and that publishing code, data, and instructions is a prerequisite to ensure experimental findings are reproducible (Buckheit and Donoho, 1995; Peng, 2011; Raff, 2019; Pineau et al., 2021).
@@
Figure 1: Example from SciCoQA, showing a specific model implementation in the paper and its implementation in the code (simplified for readability). The paper’s description and the code’s implementation mismatch, creating a paper-code discrepancy.
@@
However, the availability of code guarantees neither reproducibility nor consistency with the scientific text (as showcased in Fig. 1). In practice, implementation details can diverge from their descriptions, introducing performance variations that go unreported (Henderson et al., 2018). These discrepancies manifest in troubling ways: from “mathiness,” where equations simulate technical depth while actual gains stem from undocumented tricks (Lipton and Steinhardt, 2019), to evaluation metrics that differ in implementation (Post, 2018; Marie et al., 2021), rendering scientific comparisons invalid.
@@
Reproducibility issues and paper-code inconsistencies are typically only detected during reproduction efforts and post-publication, resulting in a waste of resources and eroding trust in science. While checklists during peer review can encourage authors to provide more reproducible code (Pineau et al., 2021; Dodge et al., 2019), ideally, reviewers check the correctness of the implementation and ensure reproducibility. However, with the rapid growth of submission numbers, reviewers are already under severe time pressure (Rogers and Augenstein, 2020) and conducting intricate code reviews is time-consuming (Collberg and Proebsting, 2016).
@@
Furthermore, the reliance on manual review is becoming increasingly impractical as science begins to scale via agentic systems that develop ideas, generate and execute code, and produce scientific articles autonomously (Boiko et al., 2023; Lu et al., 2024; Tang et al., 2025; Weng et al., 2026; Lu et al., 2026), a trend recently exemplified by peer-reviewed acceptance to an ICLR workshop (Yamada et al., 2025). While this direction has the potential to accelerate scientific discovery, it makes oversight increasingly challenging as humans cannot verify the rapidly expanding volume of output (Bowman et al., 2022). Moreover, the faithfulness of these systems is not guaranteed as LLMs suffer from limited context (Liu et al., 2024b), experience compounding errors (Dziri et al., 2023; Mirzadeh et al., 2025), and struggle with independent self-correction (Tyen et al., 2024; Wu et al., 2024). Consequently, ensuring the reliability of science at this scale demands automated tools capable of verifying that the code faithfully implements the methods reported in the scientific paper.
@@
To this end, we introduce SciCoQA, the first benchmark for detecting discrepancies between scientific papers and their code as a cross-modal verification task. Benchmarking LLMs on this task allows us to measure their potential for automated quality assurance workflows in science. To construct our dataset, we leverage issues reported in code repositories and papers dedicated to reproducing existing work, such as those from reproducibility challenges and conference tracks. While these sources are realistic and of high quality, they are sparse and limited to CS and AI. Therefore, we scale the data to various computational science domains (e.g., Physics and Quantitative Biology) by generating modifications in codebases, creating discrepancies between the paper and its implementation. Finally, we benchmark open-weight and proprietary LLMs to determine whether current systems can be deployed for detecting paper-code discrepancies. Our evaluation reveals that the best-performing models, while precise in identifying discrepancies, suffer from insufficient recall. The best models in our evaluation detect only $46.7\%$ of real-world cases, highlighting the need to develop more effective models and systems for this critical task.
@@ 2 Related Work
Recently, LLMs have been employed to detect errors in scientific papers. Early work created short CS papers with logical errors and prompted LLMs to identify the issues (Liu and Shah, 2023). However, their reliance on manually crafted papers limits scalability. Therefore, Dycke and Gurevych (2026) and Xi et al. (2025) propose dedicated pipelines that generate modifications to papers, invalidating their key claims. They evaluate whether the soundness score of generated reviews decreases or whether LLMs can spot these errors. Both find that LLMs struggle to detect most introduced errors. Besides constructing erroneous papers, recent research focuses on identifying real-world issues. Specifically, Son et al. (2025) and Zhang and Abernethy (2025) use the authors’ retraction notes from WithdrarXiv (Rao et al., 2024) and PubPeer to obtain faulty papers. While realistic, these notes are often unspecific (e.g., “Theorem 2 is incorrect”). Recently, Bianchi et al. (2025) developed a system to detect inconsistencies within the text of AI papers (e.g., incorrect calculations in tables or errors in equations) and then manually verified the detections. They found their system to be relatively precise but to struggle with recall, i.e., detecting all issues in a paper. They also noted an increasing trend of errors in scientific publications over the years, highlighting the importance of quality assurance in science.
@@
While these works share a similar spirit to ours, i.e., identifying errors in science, they are limited by solely analyzing the paper. SciCoQA extends beyond the scientific text and considers the code as well, enabling the detection of cross-modal error types, such as implementations that deviate from the paper description or omissions in the paper or the code that are crucial for understanding and reproducibility. Conceptually, our task can be viewed as a form of cross-modal claim verification (Thorne et al., 2018; Wadden et al., 2020), where the evidence source is a codebase rather than text. Furthermore, SciCoQA utilizes constructed errors to scale the quantity and diversity in the computational science domain, while also providing challenging real-world discrepancies between paper and code.
@@
Our task is also related to Code-Comment Inconsistency (CCI) detection, where natural language comments that no longer match a piece of code need to be identified, e.g., after adding a feature to a function. Early CCI research relied on rule-based systems to identify outdated comments (Ratol and Robillard, 2017). Subsequently, bi-encoders and cross-encoders have been employed (Rabbi and Siddik, 2020; Steiner and Zhang, 2022). Panthaplackel et al. (2021) further extended CCI to dynamic settings, learning to detect inconsistency arising from code updates. Recently, LLMs have been deployed for both detection and rectification (Dau et al., 2024; Rong et al., 2025). These methods deal with local and small inputs, typically taking a single function and an inline comment or docstring as input to detect the CCI. SciCoQA is substantially more ambitious: it requires performing a global alignment by reasoning over a dense scientific paper and a long, multi-file code repository to find inconsistencies.
@@
Automation in science is rapidly accelerating (Zheng et al., 2025). Early work targeted specific components of the research cycle, including ideation (Si et al., 2025; Baek et al., 2025), literature review (Wang et al., 2024; Liang et al., 2025), coding (Tian et al., 2024; Gandhi et al., 2025), and writing (Liang et al., 2024; Weng et al., 2025). A recent trend involves generating entire code repositories from a paper (Starace et al., 2025; Seo et al., 2026; Xiang et al., 2025) or automating the scientific method end-to-end, giving rise to “AI Scientists.” However, for both these tasks, validating whether the generated code faithfully reflects the paper is challenging. For example, PaperBench (Starace et al., 2025) relies on high-quality, but expensive, manual rubrics specific to each paper to verify implementation nuances. Others use generic LLM-judges without ground truth validation (Tang et al., 2025), evaluate whether the code runs (Weng et al., 2026; Xiang et al., 2025), or assign reproducibility scores based on execution outcomes (Siegel et al., 2024; Hu et al., 2025; Ye et al., 2025). Critically, these metrics can be misleading; a generated codebase may execute perfectly and achieve high performance while implementing a method that differs fundamentally from the paper’s scientific description (Beel et al., 2025). SciCoQA addresses this gap by providing a ground truth dataset of paper-code discrepancies, enabling the development of robust quality assurance models that can verify the faithfulness of (generated) papers to (generated) code. More broadly, while code benchmarks in both software engineering (Jimenez et al., 2024) and science (Tian et al., 2024) predominantly evaluate code generation, SciCoQA evaluates the complementary task of code comprehension and verification against a scientific specification. Thus, SciCoQA takes a step toward verification and reproducibility tooling essential for sustaining peer review integrity and science at scale (Wei et al., 2025b; Woodruff et al., 2026; You et al., 2026).
@@ 3 SciCoQA
The paper-code discrepancies, according to our definition, can arise from distinct authorship (where the paper writer differs from the engineer implementing the code), simplifications made in the text for readability, or code updates or experiments that were not propagated to the manuscript.
@@ 3.1 Data Collection
Fig. 2 provides an overview of our data collection. To obtain real-world instances of paper-code discrepancies, we draw from GitHub issues and reproducibility papers. Furthermore, synthetic discrepancies are generated in real scientific codebases.
@@
We first identify repositories that reference a research paper in their homepage or description field on GitHub, restricted to projects published between 2020 and 2025 and repositories with at least one issue (see §A.1 for details). [^2] The crawled repositories are not always the official implementation by the authors, but sometimes reproductions. They were published, for example, when there is no official implementation or the reimplementation uses a different framework. We still consider these repositories, as researchers might use them to base their experiments on. For each, we crawl all associated issues, yielding 1,8901{,}890 repositories with a total of 10,63610{,}636 issues. To process this large volume efficiently, we automatically classify issues with Qwen3 4B Thinking (Yang et al., 2025), prompting the model to determine whether an issue reports a paper-code discrepancy or not, resulting in $232$ candidates (prompt in §F.1). All candidates are subsequently manually filtered to ensure they meet our definition of a paper-code discrepancy, yielding $59$ discrepancies.
@@
We collect reproducibility reports from the ML Reproducibility Challenge (Pineau et al., 2019; Sinha et al., 2020, 2022, 2023), which invites participants to reproduce papers from leading ML, CV, and NLP conferences (e.g., NeurIPS, ICML, ICLR, CVPR, ICCV, ACL, EMNLP) and papers from the reproducibility tracks of SIGIR and ECIR from 2020–2025. We retain only reports that reproduce a single, open-access paper with an open-source implementation, resulting in $171$ reproducibility papers. From each report, we extract mentions of paper-code discrepancies using GPT-5, resulting in $132$ candidates (prompt in §F.3). All extracted discrepancies are manually verified to confirm adherence to our definition, yielding $65$ discrepancies.
@@
Finally, all manually filtered discrepancies are validated by Gemini 3.1 Pro and GPT-5, checking the discrepancy’s existence given the original paper, codebase, and the GitHub issue or the description extracted from the reproducibility paper. For the cases where the LLMs disagreed, we manually verified the discrepancy. During this step, the LLMs are further tasked with generating a standardized discrepancy description. We prompt the models to output 3-8 sentences and to state what the paper describes, what the code implements, and where the difference is. We use these generated descriptions as the ground truth to ensure the same format and level of verbosity for paper-code discrepancies (prompt in §F.2 and §F.4). We use two state-of-the-art LLMs to perform the validation and rephrasing to prevent potential bias in the selection and phrasing (see §C.5 for an analysis). The verification step ensures adherence to the discrepancy definition, acting as a high-precision filter. Therefore, we retain only discrepancies that can be verified given the explicit evidence, and prioritize the precision and objectivity of the ground truth. In total, we obtain $92$ discrepancies from $72$ papers, where $42$ discrepancies originate from GitHub issues and $50$ from reproducibility papers.
@@
To scale the data in size and beyond the CS/AI domain, we employ synthetic data generation. Specifically, based on our GitHub crawl, we sample repositories linked to arXiv papers and those with permissive code licenses (i.e., MIT, Apache 2.0, CC-BY, BSD), as we will be redistributing their code partially. We randomly select $102$ repositories where the paper’s arXiv subject is CS (balanced over Machine Learning (ML, cs.LG), Computer Vision (CV, cs.CV), and Natural Language Processing (NLP, cs.CL)), and $102$ for non-CS papers. [^3] We manually checked the repositories and removed codebases from our sample that were not suitable, such as repositories for survey papers or with very few files. Next, we prompt GPT-5 with the paper and code to generate five code diffs according to our discrepancy definition (prompt in §F.5). We then sample up to three of these changes, ensuring that they do not manipulate the same file and that the generated diff can be found by exact match in the original code. Finally, we add $13$ additional discrepancies that are labeled as Paper Omissions as we later find these to be most critical (see §3.2 and §5). This yields a total of $543$ discrepancies across the $204$ papers, of which $286$ discrepancies are from the CS domain, and $257$ are from other computational domains, such as Electrical Engineering and Systems Science ($114$), Physics ($51$), or Statistics ($40$).